OpenStudy (anonymous):

g(x)+sine g(x) = x^2 Find g'(0) using implicit differentiation.

OpenStudy (abb0t):

It's a bit hard to understand the question, but I think you will have: \[g'(x) + \cos(g(x)) g'(x) = 2x\] Factor out a g'(x): \[g'(x)[1+\cos(g(x))] = 2x\]

OpenStudy (abb0t):

Finally, divide: \[g'(x) = \frac{ 2x }{ 1+\cos(g(x)) }\] and plug in zero.

OpenStudy (anonymous):

So it's kinda similar to the chain rule then.

OpenStudy (abb0t):

Well, you did use chain rule. So yes. Unless you had\[\sin(x)g(x)\] in which case you would use product rule instead of chain rule.

OpenStudy (anonymous):

yup, that makes sense. Thank you!